# Fast convergence of Frank-Wolfe algorithms on polytopes

### v3

@@ 4 Facial geometry
Proof of Lemma 3.9..
Suppose that ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ with ${\mathbf{y}}\neq{\mathbf{x}}$ as otherwise (3.20) trivially holds. Consider the following modified version of $\mathfrak{v}({\mathbf{y}},{\mathbf{x}})$:
@@
Proof of Lemma 3.14..
This proof is similar, but simpler, to that of Lemma 3.9. Suppose that ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ with ${\mathbf{y}}\neq{\mathbf{x}}$ as otherwise (3.31) trivially holds. Consider the following modified version of $\mathfrak{f}({\mathbf{y}},{\mathbf{x}})$:
@@
and suppose ${\mathbf{v}},{\mathbf{u}}$ attain the value $\gamma=\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})$, that is, ${\mathbf{v}}\in F({\mathbf{y}}),{\mathbf{u}}\in F({\mathbf{x}})$ and ${\mathbf{y}}-{\mathbf{x}}=\gamma({\mathbf{v}}-{\mathbf{u}}).$ The construction of $\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})$ and the choice of $\gamma,{\mathbf{u}},{\mathbf{v}}$ imply that $F({\mathbf{u}})\cap F({\mathbf{v}})=\emptyset$. To finish, observe that $\gamma=\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})\geq\mathfrak{f}({\mathbf{y}},{\mathbf{x}})$ and thus
@@ 4.2 Lower bound on the inner and outer facial distances
Proof. The construction of $I_{F}$ and $\mathbb{I}(F)$ immediately implies that $F\subseteq\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I}{\mathbf{x}}={\mathbf{e}}_{I}\}=\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}$ for all $I\in\mathbb{I}(F)$. For the reverse inclusion, again the construction of $I_{F}$ implies that if ${\mathbf{x}}\in\mathcal{C}\setminus F$ then $\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle>e_{i}$ for some $i\in I_{F}$. Therefore $\mathcal{C}\setminus F\subseteq\mathcal{C}\setminus\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}$ and consequently $\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}\subseteq\mathcal{C}\cap F$. ∎

### v4

@@ 4 Facial geometry
Proof.
Proof of Lemma 3.9. Suppose that ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ with ${\mathbf{y}}\neq{\mathbf{x}}$ as otherwise (3.23) trivially holds. Consider the following modified version of $\mathfrak{v}({\mathbf{y}},{\mathbf{x}})$:
@@
Proof.
Proof of Lemma 3.14. This proof is similar, but simpler, to that of Lemma 3.9. Suppose that ${\mathbf{x}},{\mathbf{y}}\in\mathcal{C}$ with ${\mathbf{y}}\neq{\mathbf{x}}$ as otherwise (3.34) trivially holds. Consider the following modified version of $\mathfrak{f}({\mathbf{y}},{\mathbf{x}})$:
@@
and suppose ${\mathbf{v}},{\mathbf{u}}$ attain the value $\gamma=\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})$ in (4.2), that is, ${\mathbf{v}}\in F({\mathbf{y}}),{\mathbf{u}}\in F({\mathbf{x}})$ and ${\mathbf{y}}-{\mathbf{x}}=\gamma({\mathbf{v}}-{\mathbf{u}}).$ The construction of $\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})$ and the choice of $\gamma,{\mathbf{u}},{\mathbf{v}}$ imply that $F({\mathbf{u}})\cap F({\mathbf{v}})=\emptyset$. To finish, observe that $\gamma=\bar{\mathfrak{f}}({\mathbf{y}},{\mathbf{x}})\geq\mathfrak{f}({\mathbf{y}},{\mathbf{x}})$ and thus
@@ 4.2 Lower bound on the inner and outer facial distances
Proof. The construction of $I_{F}$ and $\mathbb{I}(F)$ immediately implies that $F\subseteq\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I}{\mathbf{x}}={\mathbf{e}}_{I}\}=\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}$ for all $I\in\mathbb{I}(F)$. For the reverse inclusion, again the construction of $I_{F}$ implies that if ${\mathbf{x}}\in\mathcal{C}\setminus F$ then $\langle{\mathbf{d}}_{i},{\mathbf{x}}\rangle>e_{i}$ for some $i\in I_{F}$. Therefore $\mathcal{C}\setminus F\subseteq\mathcal{C}\setminus\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}$ and consequently $\mathcal{C}\cap\{{\mathbf{x}}\mid D_{I_{F}}{\mathbf{x}}={\mathbf{e}}_{I_{F}}\}\subseteq\mathcal{C}\cap F=F$.∎